OpenStudy (anonymous):

Randy is driving 4 miles per hour faster than Sandy and if they are 287 miles apart after 3.5 hours then how fast is Sandy driving?

OpenStudy (amistre64):

287/4 should get you randys starting point

OpenStudy (amistre64):

287/3.5

OpenStudy (anonymous):

i set it up as 287=3.5S + 3.5R R=S+4 (R being randys speed, S being sandys speed)

OpenStudy (amistre64):

speed randy (time) + speed sandy (time) = 287

OpenStudy (amistre64):

good good

OpenStudy (anonymous):

and i got 39 for S but thats not one of my options

OpenStudy (anonymous):

idk what i did wrong

OpenStudy (amistre64):

287 = (s+4)(3.5) + 3.5s 287 = 3.5s +3.5s + 14 287-14 = 7s 273 = 7s 273/7 = s 39 = s

OpenStudy (amistre64):

maybe its a bad model :)

OpenStudy (amistre64):

randy would be going 43 and sandy 39 right?

OpenStudy (anonymous):

thats what i figured

OpenStudy (amistre64):

maybe it changes if they are starting from the same spot

OpenStudy (anonymous):

my options are 51 56 61 66 71

OpenStudy (amistre64):

randy - Sally = 287

OpenStudy (amistre64):

(3.5s + 14) - (3.5s) = 287 ?

OpenStudy (anonymous):

its 39! :(

OpenStudy (anonymous):

actually i got 0 when i did it again

OpenStudy (amistre64):

lol try randy-287 = sally

OpenStudy (amistre64):

nope

OpenStudy (anonymous):

GAHHHH my teacher sucks

OpenStudy (amistre64):

distance = speed (time) right?

OpenStudy (anonymous):

yes btw the last question you helped me with..the pools..i was looking at the wrong answer options..it wasnt 61

OpenStudy (amistre64):

ohhhh...was that the one with 7.1 radius

OpenStudy (anonymous):

yes im just gonna yell at my teacher and tell him he gave me a faulty question

OpenStudy (amistre64):

I get: Dr = Ds + 287 but dunno if its gonna work :)

OpenStudy (amistre64):

(s+4)(3.5) = s(3.5) + 287

OpenStudy (anonymous):

got 0

OpenStudy (amistre64):

if they start at different points and travel towards each other, but thats not in the question

OpenStudy (anonymous):

right they are from the same point and driving away from each other

OpenStudy (amistre64):

away from each other?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

thenDr + Ds = 287 if thats the case 7s +14 = 287 7(s+2) = 287 i still get 39

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

this class is ridiculous these are like impossible problems and i still have a whole other set of problems to start

OpenStudy (amistre64):

if sally is not moving, does that change the dynamics of our problem?

OpenStudy (anonymous):

yeah cuz then she would have no speed

OpenStudy (anonymous):

and therefore would have no distance

OpenStudy (amistre64):

think of it like this maybe....in one hour, randy if 4 miles farther.....

OpenStudy (amistre64):

randy, at the end of 3.5 hours is 14 miles farther than sally

OpenStudy (anonymous):

that wouldnt change the equation for sally's speed though..

OpenStudy (amistre64):

they have traveled a distance of 273 miles if they were traveling the same speed right?

OpenStudy (anonymous):

no because when we substitute R in for S+4 they arent

OpenStudy (anonymous):

and the question says they are travleing different speeds

OpenStudy (amistre64):

randy is 14 miles further than sally at 3.5 hours right?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

so 287 - 14 = the distance traveled by Sally in 3.5 hours right? 273 in 3.5 hours?

OpenStudy (amistre64):

speed equals distance/time

OpenStudy (amistre64):

that gives me 78 mph..... but its closer :)

OpenStudy (anonymous):

hahah and still not an option

OpenStudy (anonymous):

I am giving up for tonight, thank you for all your help

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

Gnite :)